A freezer has a coefficient of performance of 2.40. The freezer is to convert 2.10 kg of water at T = 25.0°C to 2.10 kg of ice at T = -5.0°C in one hour. (a) What amount of heat must be removed from the water at 25.0°C to convert it to ice at -5.0°C? (b) How much electrical energy is consumed by the freezer during this hour? c) How much wasted heat is delivered to the room in which the freezer sits?

Please help me with this problem but please only help if you know how to do it 100% and please circle answers
Answer #1

(1)

Let Q the amount of heat to be removed,

Q = m [C(0-25) - 79.8 + C'(-5 -0)]

Q = - 2100*(25 +79.8 + 0.50*5) = - 225330 cal = -942780.72 J

(2)

The electrical energy consumed is

E = |Q|/COP = 942780.72/2.40 = 392825.3 J

(3)

The heat rejected to the room is:

Q' = |Q| + E = 1335606.02 J